9. A DNA molecule has 180 base pairs and is 20 percent adenine. How many cytosine nucleotides are present in this molecule of DNA?
Answer #1

Answer:

108 molecules of cytosine are present in this molecules of DNA.

Explanation:

Each base pair consists of two bases.

180 base pairs = 180*2= 360 bases

20% adenines means= 20 * 360/100= 72 adenine molecules

A=T

72 Adenines = 72 Thymines

360-144 = 216

216 = G + C

G = 108

C= 108